The mole fraction of nickel(II) acetate, Ni(CH3COO)2, in an aqueous solution is 3.76x102 The percent by mass of nickel(II) acetate in the solution is 0 Submit AnswerAn aqueous solution is 12.0% by mass ammonia NH3 The mole fraction of ammonia in the solution is Submit AnswerUse the References to access important values if needed for this question. EPA regulations for harmful contaminants in drinking water set an upper limit of 4.00x103 mg/L for beryllium. Suppose that your water supply contained Be at a level of 0.516 ppb. (a) What is this concentration expressed in mg/L? mg/L (b) Would your water supply pass EPA regulations? Submit AnswerUse the References to access important values if needed for this question. The EPAs secondary standards for contaminants that may cause cosmetic or aesthetic effects in drinking water suggest an upper limit of 5.00x102 mg/L for manganese. If 1.30x104 liters of water in a storage tank contains 0.172 grams of Mn, what is the contaminant level in ppm? Is this level acceptable based on EPA guidelines? ppm Submit Answer
Answer #1

1)

Let the total mol in solution be 1 mol

Then mol of Ni(CH3COO)2 is 3.76*10^-2 mol

mol of H2O = 1 - 3.76*10^-2 mol

= 0.9624 mol

Molar mass of Ni(CH3COO)2,

MM = 1*MM(Ni) + 4*MM(C) + 6*MM(H) + 4*MM(O)

= 1*58.69 + 4*12.01 + 6*1.008 + 4*16.0

= 176.778 g/mol

use:

mass of Ni(CH3COO)2,

m = number of mol * molar mass

= 3.76*10^-2 mol * 1.768*10^2 g/mol

= 6.647 g

Molar mass of H2O,

MM = 2*MM(H) + 1*MM(O)

= 2*1.008 + 1*16.0

= 18.016 g/mol

use:

mass of H2O,

m = number of mol * molar mass

= 0.9624 mol * 18.02 g/mol

= 17.34 g

mass % = mass of Ni(CH3COO)2 * 100 / total mass

= 6.647 * 100 / (6.647 + 17.34)

= 27.7 %

Answer: 27.7 %

Only 1 question at a time please